OpenStudy (anonymous):

A geometric sequence is defined by the explicit formula an = 5(-4)n-1. What is the recursive formula for the nth term of this sequence? A. an = -5an-1 B. an+1 = 4an C. an+1 = 5an D. an = -4an-1 someone please help me...

OpenStudy (solomonzelman):

\(\large\color{slate}{ a_n = 5(-4)^n-1 }\) like this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

except its n-1

OpenStudy (solomonzelman):

oh, \(\large\color{slate}{ a_n = 5(-4)^{n-1} }\) this ?

OpenStudy (anonymous):

yes!!

OpenStudy (solomonzelman):

\(\large\color{slate}{ a_n = 5(-4)^{n-1} }\) \(\large\color{slate}{ a_n = 5(-4)^{n} (-4)^{-1} }\) \(\large\color{slate}{ \displaystyle a_n = 5(-4)^{n} \frac{1}{4} }\) \(\large\color{slate}{ \displaystyle a_n = \frac{5}{4} (-4)^{n} }\)

OpenStudy (solomonzelman):

that is not any of the options though, is it....

OpenStudy (solomonzelman):

your options are very difficult to read, perhaps you can screenshot them or draw?

OpenStudy (solomonzelman):

I made a mistake in my conslusion

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle a_n = -\frac{5}{4} (-4)^{n} }\) because it is negative 1/4 and on...... but will try to get it to one of your options

OpenStudy (anonymous):